Consider a mixture of 2 mol of helium and 4 mol of oxygen. Compute the speed of sound in this gas mixture at 300 K

1
Expert's answer
2022-05-22T14:38:45-0400

"v=\sqrt{\frac{\gamma RT}M}=380~\frac ms."